# Removal and Selection: Improving RGB-Infrared Object Detection via Coarse-to-Fine Fusion

### v4

@@ Abstract
Object detection in visible (RGB) and infrared (IR) images has been widely applied in recent years. Leveraging the complementary characteristics of RGB and IR images, the object detector provides reliable and robust object localization from day to night. Most existing fusion strategies directly input RGB and IR images into deep neural networks, leading to inferior detection performance. However, the RGB and IR features have modality-specific noise, these strategies will exacerbate the fused features along with the propagation. Inspired by the mechanism of the human brain processing multimodal information, in this paper, we introduce a new coarse-to-fine perspective to purify and fuse two modality features. Specifically, following this perspective, we design a Redundant Spectrum Removal module to coarsely remove interfering information within each modality and a Dynamic Feature Selection module to finely select the desired features for feature fusion. To verify the effectiveness of the coarse-to-fine fusion strategy, we construct a new object detector called the Removal and Selection Detector (RSDet). Extensive experiments on three RGB-IR object detection datasets verify the superior performance of our method. The source code and results are available at https://github.com/Zhao-Tian-yi/RSDet.git
@@
RGB-Infrared Object Detection, Coarse-to-Fine Fusion Strategy, Multisensory Fusion, Mixture of Experts.
@@ I Introduction
Object detection on single modality is one of the fundamental tasks in computer vision, garnering significant attention and finding applications across diverse fields such as surveillance[1, 2], remote sensing[3, 4, 5], and autonomous driving[6, 7]. However, object detection methods reliant solely on one single modality have shown susceptibility to various challenges [8], like illumination, weather, adversarial attack, occlusion, etc. Therefore, the multiple sensor modalities fusion method has come into view[9, 10, 11, 12, 13]. With the development of various sensors, multimodal images are collected and applied in more and more application fields. Among them, visible (RGB) and infrared (IR) sensors are widely utilized due to their complementary imaging characteristics. Specifically, visible images can reflect rich texture and color information of the object but are greatly affected by low-light environments. In contrast, infrared images can reflect the thermal information of the object and can clearly reflect the outline of the object under poor illumination conditions, but lack the texture details of the object. In recent years, researchers have focused on the visible and infrared feature-level fusion  [14, 15, 16, 11], which can help achieve better performance on downstream tasks (e.g., object detection).
@@
In RGB-IR object detection, the effective feature fusion method of RGB and IR images is crucial. Most existing RGB-IR object detection methods extract the modality-specific features from RGB and IR images independently, and then directly perform addition or concatenation operations on these features [17, 18, 19], as shown in Figure 1(a). Without explicit cross-modal fusion, the “Late fusion” strategy is therefore limited in learning the complementary information, resulting in inferior performance. To further explore the optimal fusion strategies, many researchers have considered the “Halfway fusion” strategy to design an interaction module between different modality features [20, 21, 22], as shown in Figure 1(b). For instance, Zhou et al. [20] construct the DMAF module to tap the difference between RGB and IR modalities which brings more useful information at the feature level. Xie et al. [21] introduce a novel dynamic cross-modal module that aggregates local and global information from RGB and IR modalities, etc. Although these methods have achieved encouraging improvements, they ignore the negative impact of the modality-noise along with the propagation, which would fail to achieve complementary fusion.
Actually, when confronted with multi-modal information, our brains initially establish rules to filter out interfering information and then meticulously select the desired information, a process that has been modeled in cognitive theory (“Attenuation Theory” [23]). This approach can be likened to a coarse-to-fine process, inspiring us to introduce a fresh perspective for fusing RGB and IR features. As shown in Figure 1(c), we design a new fusion strategy called “Coarse-to-Fine Fusion” to achieve complementary feature fusion.
“Coarse” indicates that our method begins with the filter out the interfering information and thus can coarsely remove the irrelevant spectrum. To this end, since the redundant information in an image also exists in its frequency spectrum [24], we propose a Redundant Spectrum Removal (RSR) module to filter out coarsely in the frequency domain. Specifically, we convert each source image into frequency space and introduce a dynamic filter to adaptively reduce irrelevant spectrum within RGB and IR modalities. As for “Fine”, it indicates that our fusion strategy conducts finely select features after the coarse removal. We design a Dynamic Feature Selection (DFS) module to meticulously select the desired features between RGB and IR modalities. In this module, we weighted different scale features required for object detection by exploring a mixture of scale-aware experts. Figure 2 gives an example of our Coarse-to-Fine fusion strategy. To evaluate the effectiveness of the coarse-to-fine strategy, we build a novel framework for RGB-IR object detection called Removal and Selection Detector (RSDet).
@@
  • Following the coarse-to-fine fusion strategy, we propose a Redundant Spectrum Removal module, which introduces a dynamic spectrum filter to adaptively reduce irrelevant information in the frequency domain.
  • We also design a Dynamic Feature Selection module, which utilizes a mixture of scale-aware experts to weigh different scale features for the RGB-IR feature fusion.
@@ II-A RGB-IR Object Detection
In recent years, thanks to the development of deep learning technology and several visible and infrared datasets being proposed [25, 26, 27], the RGB-IR object detection (also known as multispectral object detection) task has gradually attracted more and more attention. To fully explore the effective information between visible and infrared images, some researchers focus on the complementarity between the two modalities starting from the illumination conditions. Guan et al. [28] and Li et al. [29] first propose the illumination-aware modules to allow the object detectors to adjust the fusion weight based on the predicted illumination conditions. Moreover, Zhou  et al. [20] analyze and address the modality imbalance problems by designing two feature fusion modules called DMAF and IAFA. Recently, an MSR memory module [30] was introduced to enhance the visual representation of the single modality features by recalling the RGB-IR modality features, which enables the detector to encode more discriminative features. Yuan et al. [31] propose a transformer-based RGB-IR object detector to further improve the object detector performance. Different from these works, We are inspired by the classic theory of cognitive psychology, Attenuation Theory, to emulate the way the human brain processes information from multiple sources. We propose a coarse-to-fine fusion strategy to fuse RGB and IR features for the RGB-IR object detection task.
@@ II-C Mixture of Experts
The Mixture-of-Experts (MoE) model [36, 37] has demonstrated the ability to dynamically adapt its structure based on varying input conditions. Several studies have been dedicated to the theoretical exploration of MoE [38, 39, 40, 41], focusing on the sparsity, training effectiveness, router mechanisms, enhancing model quality, etc. Besides, some researchers also concentrated on leveraging the MoE model for specific downstream tasks. For example, Gross et al. [42] observed that a hard mixture-of-experts model can be efficiently trained to good effect on large-scale multilabel prediction tasks. Cao et al. [43] proposed a mixture of local-to-global experts (MoE-Fusion) mechanisms by integrating MoE structure into image fusion tasks. Chen et al. [44] addressed the multi-task learning challenge by implementing a cooperative and specialized mechanism among experts. In our proposed method, the MoE model is integrated into the RGB-IR object detection task. Specifically, we establish multi-scale experts to gate modality features at different scales and leverage its dynamic fusion mechanism to facilitate complementary fusion.
III Method
III-A Coarse-to-Fine Fusion
The inspiration for the proposed “Coarse-to-Fine” Fusion strategy stems from the cognitive models of human processing of multiple sources of information, named Selective Attention Theory in cognitive psychology (e.g., “Broadbent’s filter model” [45] and “reisman’s Attenuation Model” [23]), which is one of the cornerstones of attention mechanism theory in cognitive psychology. The Selective Attention Theory posits that when the human brain processes multiple external stimuli, it first filters out or attenuates some unimportant or even useless information. Then, it further processes the information in a more refined manner before storing it in working memory, as illustrated in Figure 3. The Attenuating Filter corresponds to the Redundant Spectrum removal module (RSR), which effectively filters out unimportant or interfering information based on some properties and rules when confronted with multi-sensory information. The Hierarchy of Analyzers corresponds to our proposed Dynamic Feature module (DFS), engaging in meticulous, higher-level analysis and processing to extract valuable insights. Finally, the processed information enters the brain’s working memory, corresponding to the output of the fused feature by our “Coarse-to-Fine Fusion” strategy and fed into the detection network.
@@ III-A Coarse-to-Fine Fusion
According to the above inspiration, our “Coarse-to-Fine” Fusion strategy is designed as shown in Figure 4, “Coarse” corresponding to the Redundant Spectrum Removal (RSR) module to filter out coarsely in the frequency domain and ‘Fine’ corresponding to the Dynamic Feature Selection (DFS) module to meticulously select the desired features between RGB-IR modalities. Since the two modality features are usually intersection, we introduce disentangled representation learning [33] to purify and decouple them for complementary fusion. Thus, we integrate the above two modules into the shared-specific structures to implement our coarse-to-fine fusion. Specifically, we first perform the RSR module to remove interfering information. Then, we introduce the shared-specific structure to extract the two modality-specific multiscale features $C_{\text{I-spe}}$ and $C_{\text{V-spe}}$, which uses ResNet50 [46] as the backbone network. As for the shared features $C_{\text{sha}}$, we also employ several Resblocks as the feature extractor. After that, these different scale features $C_{\text{I-spe}}$ and $C_{\text{V-spe}}$ are dynamically aggregated by the proposed mixture of scale-aware experts and obtain the final specific feature $C_{\text{spe}}$ in the DFS module. Finally, the specific feature $C_{\text{spe}}$ and the shared feature $C_{\text{sha}}$ are added together for the final object detection task, which can be expressed as:
@@
Figure 3: Illustration of Treisman’s Attenuation Model
Figure 4: The overall architecture of RSDet. The entire network follows the “Coarse-to-Fine” fusion and comprises a two-stream backbone network taking RGB and IR inputs, a Redundant Spectrum Removal (RSR) module, and a Dynamic Feature Selection (DFS) module.
@@ III-B Redundant Spectrum Removal
For the “Coarse Removal”, since the frequency domain obeys the nature of global modeling, we propose a Redundant Spectrum Removal (RSR) module to filter out coarsely in the frequency domain. we first transform each source image into the frequency domain and predict a dynamic filter to adaptively reduce irrelevant spectrum within RGB and IR modalities. Specifically, the paired RGB image $V\in\mathbb{R}^{H\times W\times 3}$ and IR image $I\in\mathbb{R}^{H\times W\times 1}$ are token as the RSR module input. They are first subjected to the Discrete Fourier Transform ($\operatorname{DFT}(\cdot)$) and get the frequency domain image $F_{I}(u,v)$ and $F_{V}(u,v)$:
$F_{I}(u,v)=\operatorname{DFT}(I),~{}F_{V}(u,v)=\operatorname{DFT}(V).$ (2)
The filter prediction network is conducted to dynamically generate the redundant spectrum filter, which is illustrated in Figure 4. For each modality image, we perform an embedding operation on the entire image get a feature vector, and reshape it to match the size of the original image:
${M}_{l_{I}}=\operatorname{Embed}(I),$ ${M}_{l_{V}}=\operatorname{Embed}(V),$ (3)
$M_{m_{I}}=\operatorname{Reshape}(M_{l_{I}}),$ $M_{m_{V}}=\operatorname{Reshape}(M_{l_{V}}).$
Then, for adaptively reducing irrelevant spectrum according to different inputs by the amplitude information of the original image, we utilize the $\operatorname{top}K$ operation on $M_{m}$. In specific, the positions don’t correspond to the $\operatorname{top}K$ values are set to 0:
$H_{I}(u,v)=\operatorname{top}K(M_{m_{I}}),H_{V}(u,v)=\operatorname{top}K(M_{m_{V}}).$ (4)
we apply $H(u,v)$ to filter the frequency domain image $F(u,v)$ for reducing irrelevant spectrum:
$F_{I}^{{}^{\prime}}(u,v)=F_{I}(u,v)\otimes H_{I}(u,v)$ (5)
@@
Finally, the filtered $F_{I}^{{}^{\prime}}(u,v)$ and $F_{V}^{{}^{\prime}}(u,v)$ are subjected to the Inverse Discrete Fourier Transform $\operatorname{IDFT}(\cdot)$ to transform the image back to the spatial domain. Thus, we obtain the final irrelevant spectrum removed images $I^{{}^{\prime}}$ and $V^{{}^{\prime}}$.

### v6

# Removal then Selection: A Coarse-to-Fine Fusion Perspective for RGB-Infrared Object Detection

@@ Abstract
In recent years, object detection utilizing both visible (RGB) and thermal infrared (IR) imagery has garnered extensive attention and has been widely implemented across a diverse array of fields. By leveraging the complementary properties between RGB and IR images, the object detection task can achieve reliable and robust object localization across a variety of lighting conditions, from daytime to nighttime environments. Most existing multi-modal object detection methods directly input the RGB and IR images into deep neural networks, resulting in inferior detection performance. We believe that this issue arises not only from the challenges associated with effectively integrating multimodal information but also from the presence of redundant features in both the RGB and IR modalities. The redundant information of each modality will exacerbates the fusion imprecision problems during propagation. To address this issue, we draw inspiration from the human brain’s mechanism for processing multimodal information and propose a novel coarse-to-fine perspective to purify and fuse features from both modalities. Specifically, following this perspective, we design a Redundant Spectrum Removal module to remove interfering information within each modality coarsely and a Dynamic Feature Selection module to finely select the desired features for feature fusion. To verify the effectiveness of the coarse-to-fine fusion strategy, we construct a new object detector called the Removal then Selection Detector (RSDet). Extensive experiments on three RGB-IR object detection datasets verify the superior performance of our method. The source code and results are available at https://github.com/Zhao-Tian-yi/RSDet.git
@@
RGB-Infrared Object Detection, Coarse-to-Fine Fusion, Multisensory Fusion, Mixture of Scale-aware Experts.
@@ I Introduction
Object detection is one of the fundamental tasks in computer vision, attracting substantial attention and finding applications in a wide range of fields such as surveillance [1], remote sensing [2, 3, 4], autonomous driving [5, 6], etc. However, relying solely on visible imagery for object detection has been shown to be susceptible to various challenges [7], like limited illumination, similar appearance of background and foreground, adversarial attack, etc. With the development of sensor technology, various modality images are collected and applied in more and more application fields Therefore, the multi-modal fusion methods [8, 9, 10, 11] have come into view. Among them, visible (RGB) and infrared (IR) sensors are widely utilized due to their complementary imaging characteristics. Specifically, IR images can clearly provide the outline of the object under poor illumination conditions through the temperature-attached thermal radiation information of the object, which can be regarded as complementary information for RGB images. Thus, in recent years, researchers have focused on the RGB and IR feature-level fusion  [12, 13], which can help achieve better performance on downstream tasks (e.g., object detection).
@@
In RGB-IR object detection, an effective feature fusion method of RGB and IR images is crucial. Most existing RGB-IR object detection methods extract the modality-specific features from RGB and IR images independently, and then directly perform addition or concatenation operations on these features [14, 15, 16], as shown in Figure 1(a). Without explicit cross-modal fusion, the “Late fusion” strategy is therefore limited in learning the complementary information, resulting in inferior performance. To further explore the optimal fusion strategies, many researchers have explored the “Halfway fusion” strategy to design an interaction module between different modality features [17, 18, 19], as shown in Figure 1(b). For instance, Zhou et al. [17] construct the MBNet to tap the difference between RGB and IR modalities which brings more useful information at the feature level. Xie et al. [18] introduce a novel dynamic cross-modal module that aggregates local and global features from RGB and IR modalities, etc. Although these methods have achieved encouraging improvements, they explicitly enforce the complementary information learning and ignore the negative impact of redundancy features along with the propagation, which would difficult to achieve complementary fusion.
Actually, when confronted with multi-modal information, our brains initially establish rules to filter out interfering information and then meticulously select the desired information, a process that has been modeled in cognitive theory (“Attenuation Theory” [20]). This theory can be likened to a coarse-to-fine process, inspiring us to introduce a new perspective for fusing RGB and IR features. As shown in Figure 1(c), we design a new fusion strategy called “Coarse-to-Fine Fusion” to achieve complementary feature fusion. “Coarse” indicates that our method begins with the filter out the interfering information and thus can coarsely remove the irrelevant spectrum. To this end, since the redundant information in an image also exists in its frequency spectrum [21], we propose a Redundant Spectrum Removal (RSR) module to filter out coarsely in the frequency domain. Specifically, we convert each source image into frequency space and introduce a dynamic filter to adaptively reduce irrelevant spectrum within RGB and IR modalities. As for “Fine”, it indicates that our fusion strategy conducts finely select features after the coarse removal. We design a Dynamic Feature Selection (DFS) module to meticulously select the desired features between RGB and IR modalities. Therefore, we can weight different scale features required for object detection by exploring a mixture of scale-aware experts. Figure 2 visualizes an example results of our Coarse-to-Fine fusion strategy. To evaluate the effectiveness of the coarse-to-fine strategy, we construct a novel framework that embeds our coarse-to-fine fusion for RGB-IR object detection called Removal then Selection Detector (RSDet).
@@
  • Following the coarse-to-fine fusion perspective, we propose a Redundant Spectrum Removal module, which introduces a dynamic spectrum filter to adaptively reduce irrelevant information in the frequency domain. We also design a Dynamic Feature Selection module, which utilizes a mixture of scale-aware experts to weigh different scale features for the RGB-IR feature fusion.
@@
The rest of this paper is organized as follows: In Section II, we briefly introduce an overview of relevant studies in RGB-IR Object detection, Shared-Specific Representation learning, and Mixture of Experts. The details of our proposed method are discussed in Section III and several experiments are conducted to validate the proposed model in Section IV. Finally, in Section V, a conclusion is made for this paper.
Figure 3: The overall framework of coarse-to-fine fusion strategy, which mainly consists of the Redundant Spectrum Removal and the Dynamic Feature Selection module. Based on this fusion strategy, a complete object detector named Remove and Select Detector (RSDet) is constructed to evaluate its effectiveness.
@@ II-A RGB-IR Object Detection
In recent years, thanks to the development of deep learning technology and several visible and infrared datasets being proposed [22, 23], the RGB-IR object detection (also known as multispectral object detection) task has gradually attracted more and more attention. To fully explore the effective information between visible and infrared images, some researchers focus on the complementarity between the two modalities starting from the illumination conditions. Guan et al. [24] and Li et al. [25] first propose the illumination-aware modules to allow the object detectors to adjust the fusion weight based on the predicted illumination conditions. Moreover, Zhou  et al. [17] analyze and address the modality imbalance problems by designing two feature fusion modules called DMAF and IAFA. Recently, an MSR memory module [26] was introduced to enhance the visual representation of the single modality features by recalling the RGB-IR modality features, which enables the detector to encode more discriminative features. Yuan et al. [23] propose a transformer-based RGB-IR object detector to further improve the object detector performance.
In this paper, we draw inspiration from the Attenuation Theory [20] to emulate how the human brain processes information from multiple sources and propose a coarse-to-fine fusion perspective to utilize RGB and IR features for object detection, which enables the fused features to be more discriminative. Meanwhile, our design of the fusion module ensures the integration effect while also guaranteeing lightweight compared to other fusion modules, as well as lower computational consumption.
@@ II-C Mixture of Experts
The Mixture-of-Experts (MoE) model [31, 32] has demonstrated the ability to dynamically adapt its structure based on varying input conditions. Several studies have been dedicated to the theoretical exploration of MoE [33], focusing on the sparsity, training effectiveness, router mechanisms, enhancing model quality, etc. Besides, some researchers also concentrated on leveraging the MoE model for specific downstream tasks. For example, Gross et al. [34] observed that a hard mixture-of-experts model can be efficiently trained to good effect on large-scale multilabel prediction tasks. Cao et al. [35] proposed a mixture of local-to-global experts (MoE-Fusion) mechanisms by integrating MoE structure into image fusion tasks. Chen et al. [36] addressed the multi-task learning by implementing a cooperative and specialized mechanism among experts.
In our proposed method, we introduce the MoE model into the RGB-IR object detection task and propose a mixture of scale-aware experts. Specifically, we design multi-scale experts and leverage its dynamic fusion mechanism to select the desired scale-specific features.
Figure 4: Illustration of Treisman’s Attenuation Model [20]
III The Proposed Method
III-A “Coarse-to-Fine” Fusion
The proposed “Coarse-to-Fine” Fusion strategy is inspired by cognitive models of human information processing, specifically the Selective Attention Theory in cognitive psychology. Notable examples include Broadbent’s Filter Model [37] and Treisman’s Attenuation Model [20]. These models serve as a cornerstone of attention mechanism theory in cognitive psychology. As shown in Figure 4, Treisman’s Attenuation Model posits that when the human brain processes multiple stimuli, it first attenuates unimportant or irrelevant messages based on specific criteria. Then it processes the remaining messages in a more refined manner which conducts detailed, hierarchical analysis and processing to extract meaningful features and insights. Finally, the processed messages enter the brain’s working memory.
@@ III The Proposed Method
Inspired by Treisman’s Attenuation Model, we design the “Coarse-to-Fine” Fusion strategy. “Coarse” corresponds to the Redundant Spectrum Removal (RSR) module to filter out coarsely in the frequency domain, and ‘Fine’ corresponds to the Dynamic Feature Selection (DFS) module to meticulously select the desired features between RGB-IR modalities. Since the two modality features often intersect, we introduce disentangled representation learning [28] to purify and decouple them for complementary fusion. As shown in Figure 3, we integrate the RSR and DFS modules into shared-specific structures to implement the “Coarse-to-Fine” Fusion. Firstly, input the RGB ($V$) and IR ($I$) images to the RSR module separately, removing interfering information to obtain the images $V^{{}^{\prime}}$ and $I^{{}^{\prime}}$ with the irrelevant redundant spectra removed. Then, we introduce the shared-specific structure to extract the two modality-specific multi-scale features $C_{\text{I-spe}}$ and $C_{\text{V-spe}}$, which uses ResNet as the backbone network. As for the shared features $C_{\text{sha}}$, we also employ several Resblocks as the feature extractor. After that, these different scale features $C_{\text{I-spe}}$ and $C_{\text{V-spe}}$ are input to the DFS module, which can dynamically aggregate them by the proposed mixture of scale-aware experts and obtain the specific feature $C_{\text{spe}}$. Finally, the specific feature $C_{\text{spe}}$ and the shared feature $C_{\text{sha}}$ are added together to get the final fused feature $C$, which can be expressed as:
@@ III-B Redundant Spectrum Removal
For ‘Coarse’, we choose to process the image in the frequency domain, due to the frequency domain having inherent global modeling properties, and only through positional multiplication operations can filter out features of the same frequency band in the entire image. But, it is difficult to handle the tight coupling of object features in the spatial domain. Therefore, we propose a Redundant Spectrum Removal (RSR) module to perform coarse filtering in the frequency domain. We first transform each input image into the frequency domain and predict a dynamic filter to attenuate irrelevant spectrum within RGB and IR modalities adaptively.
Figure 5: Example visualization of the learnable filter $H_{I}(u,v)$ and $H_{V}(u,v)$.
Specifically, the paired RGB image $V\in\mathbb{R}^{H\times W\times 3}$ and IR image $I\in\mathbb{R}^{H\times W\times 1}$ are taken as the RSR module input. They are subjected to the Discrete Fourier Transform ($\operatorname{DFT}(\cdot)$) and get the frequency domain image $F_{I}(u,v)$ and $F_{V}(u,v)$:
$F_{I}(u,v)=\operatorname{DFT}(I),$ (2)
$F_{V}(u,v)=\operatorname{DFT}(V).$
The filter prediction network is designed to dynamically generate the redundant spectrum filter based on the amplitude information $|F_{I}(u,v)|$ and $|F_{V}(u,v)|$ of the original images, which is illustrated in Figure 3. For each modality image, we perform a simple encoder on the amplitude image to obtain a feature embedding:
${M}_{l_{I}}=\operatorname{Encoder}_{I}(|F_{I}(u,v)|),$ (3)
${M}_{l_{V}}=\operatorname{Encoder}_{V}(|F_{V}(u,v)|).$
Each value of embedding ${M}_{l_{I}},{M}_{l_{V}}$ $\in\mathbb{R}^{m}$ represents the importance of different patch regions of $F_{I}(u,v)$ and $F_{V}(u,v)$ image. Then to fully retain the effective spectrum components while attenuating the useless ones, we utilize the $\operatorname{top}K$ operation on $M_{l_{I}}$ and $M_{l_{V}}$:
$M_{m_{I}}=\operatorname{top}K(M_{l_{I}}),$ (4)
$M_{m_{V}}=\operatorname{top}K(M_{l_{V}}).$
Next, we use nearest neighbor interpolation to reshape the embedding to match the size of the original image, obtaining the filters $H_{I}(u,v)$ and $H_{V}(u,v)$:
$H_{I}(u,v)=\operatorname{Reshape}(M_{m_{I}}),$ (5)
$H_{V}(u,v)=\operatorname{Reshape}(M_{m_{V}}).$
Intuitively, we visualize the learnable filters $H_{I}(u,v)$ and $H_{V}(u,v)$ as shown in Figure 5. From the ’Filtered Amplitude’, we can observe that the learned filters remove some high-frequency noise in each modality, which can be considered as redundant information for the object detection task.
Subsequently, we perform element-wise multiplication of $H_{I}(u,v)$ and $H_{V}(u,v)$ with the frequency domain images $F_{I}(u,v)$ and $F_{V}(u,v)$:
$F_{I}^{{}^{\prime}}(u,v)=F_{I}(u,v)\otimes H_{I}(u,v),$ (6)
@@
Finally, the filtered $F_{I}^{{}^{\prime}}(u,v)$ and $F_{V}^{{}^{\prime}}(u,v)$ are subjected to the Inverse Discrete Fourier Transform, denoted as $\operatorname{IDFT}(\cdot)$, to transform the images back to the spatial domain. This process yields the images $I^{{}^{\prime}}$ and $V^{{}^{\prime}}$, with the redundant and irrelevant spectrum removed.
@@
Figure 6: Visualization of the features learned by each expert ($C_{I-Spe}$ and $C_{V-Spe}$) and the fused feature ($C_{Spe}$). To facilitate a clearer observation, we use red boxes to highlight the objects and overlay the features onto the original RGB or IR image. From left to right the feature scale from large to small, corresponding to the objects size from small to large.